What is the measurement in pounds to fill a cylinder container with a diameter of 5 in. and a length of 4 feet?
Answer by checkley71(8403)

SORRY: THIS IS THE AREA NOT THE VOLUME VALUE.
-----------------------------------------------------------------
AREA=PI*R^2=3.14*2.5^2=3.14*6.25=15.7 SQ INCHES FOR THE TOP AREA
4 FEET=48 INCHES
15.7*48=753.6 ANSWER IN CUBIC INCHES.
-----------------------------------------------------------
IF YOU WANT CUBIC FEET THEN DIIVIDE BY 12*12*12=1728
753.6/1728=.436 ANSWER IN CUBIC FEET.
----------------------------------------------------------
STILL NEED TO KNOW WHAT YOU'RE FILLING THE PIPE WITH!!!!!!!
-----------------------------------------------------------
NEED TO KNOW WAHT YOU ARE FILLING THIS PIPE WITH IN ORDER TO FIND THE WEIGHT.